OpenStudy (moongazer):

given that kx +6y = 1 and 2x + 3y = 1, what should be the value of k so that there will be no solution? Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (aroub):

is this question about slope ?

OpenStudy (anonymous):

At first glance it would seem that k = 4

OpenStudy (anonymous):

Put this into the equation then try to solve simultaneously 4x + 6y = 1 2x + 3y = 1 multiply equation 2 by 2 4x + 6y = 1 4x + 6y = 2 Subtract equation 1 from equation 2 Gives an answer 0=1 So k = 4

OpenStudy (anonymous):

Does that make sense?

OpenStudy (moongazer):

where did you get 4?

OpenStudy (anonymous):

I wanted to equate the LHS of both equations

OpenStudy (anonymous):

...and I could see that 3y could easily become 6y by multiplication hence 2x can become 4x

OpenStudy (anonymous):

thats a correct solution gian theres his medal moongazer?

OpenStudy (anonymous):

Thank you Jimmy!